OpenStudy (anonymous):

x^2-6x+b=y standard form please

OpenStudy (anonymous):

what do you ean by standart form?

OpenStudy (anonymous):

mean*

OpenStudy (anonymous):

i need the y on the other side

OpenStudy (anonymous):

that's impossible.... you have one equation for two unknowns...

OpenStudy (anonymous):

is there any more info?

OpenStudy (anonymous):

Find the vertex of the parabola by completing the square.

OpenStudy (anonymous):

ohh, :)

OpenStudy (anonymous):

x2-6x +9 = (x-3)2 so (x-3)2 +b -9=y vertex(3,b-9)

OpenStudy (anonymous):

not an option

OpenStudy (anonymous):

A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

its not d

OpenStudy (anonymous):

any info about b?

OpenStudy (anonymous):

my answer is correct, just depends on what b is

OpenStudy (anonymous):

i just no its not d cause it was wrong

OpenStudy (anonymous):

3,–1)

OpenStudy (anonymous):

the only answer with first coordenate equal to 3 is B, so that should be

OpenStudy (anonymous):

it means b =8

OpenStudy (anonymous):

ok